A business analyst is hired by a local real estate firm to forecast how condo sales on a quarterly basis over the next several years. The firm provides the analyst with three years worth of quarterly condo sales data. Noticing that the data exhibits increasing trend and a seasonal pattern, the analyst chooses to create a multiple regression model for generating the forecasts:

y = 120 + 22 Qtr1 + (-70) Qtr2 + (-120) Qtr3 + 5 t

In the equation above, Qtr1, Qtr2, and Qtr3 are dummy variables corresponding to the first, second, and third quarters of the calendar year and t is the time period number, starting with t = 1 for the first quarter of year 1.

Using the model described above, what is the forecast for the number of condo sales the firm can expect during the second quarter of year 5?

(Hint: the second quarter of year 5 would be time period t = 18.)
Answer #1

necessary data:-

t = (4*4)+2 = 18

[as 1 year = 4 quarters...we are calculating second quarter of year 5 ]

Qtr1, Qtr2, and Qtr3 are dummy variables(variables that take two values, 0 and 1)..so, here, we will use Qtr2= 1 (as we are interested in second quarter of year 5 ) and Qtr1=0 , Qtr3 =0

the multiple regression equation is:-

y = 120 + 22 Qtr1 + (-70) Qtr2 + (-120) Qtr3 + 5 t

the forecast for the number of condo sales the firm can expect during the second quarter of year 5 is:-

= 120 + (22*0) + (-70*1) + (-120*0) + (5*18)

= 140

*** if you have any doubt regarding the problem please write it in the comment box.if you are satisfied please give me a LIKE if possible...